Please help me solve this question.
SOlve each equation. Chech for extraneous solutions.
2|4y+1|=4y+10

Answer by lwsshak3(11628)

SOlve each equation. Chech for extraneous solutions.
2|4y+1|=4y+10
**
Two conditions may satisfy this expression:
2|4y+1|=4y+10
8y+2=4y+10
4y=8
y=2
or
2*-(4y+1)=4y+10
-8y-2=4y+10
-12y=12
y=-1
..
solutions: y=2 and -1
..
Check:
for y=2
2|4y+1|=4y+10
2*9=8+110
18=18
..
for y=-1
2|4y+1|=4y+10
2*|-3|=-4+10
6=6
no extraneous solutions